# Experimental End-to-End Optimization of Directly Modulated Laser-based IM/DD Transmission

### v1

@@ Abstract
Index Terms:
Directly modulated laser, end-to-end learning, intensity modulation direct detection, nonlinearity, chromatic dispersion, chirp.
@@ I Introduction
The explosive growth of internet services and data traffic, driven by the surge in cloud computing and artificial intelligence (AI) applications, is placing considerable pressure on the global communications infrastructure [1]. Specifically, short-reach intra-data center interconnects must continue to increase data throughput while maintaining low costs. Intensity modulation and direct detection (IM/DD) is the dominating technology for such interconnects [2, 3, 4]. Directly modulated lasers (DMLs) are an attractive technology for short-reach IM/DD optical links, offering low energy consumption, compact size, and affordability [5, 6, 7, 8, 9, 10]. This makes them well-suited for communication systems demanding high data rates at minimal cost. DMLs are simpler and potentially more efficient than alternative transmitter technologies [9]. However, the limited modulation bandwidth of DMLs can restrict their achievable symbol rates [11, 12, 13]. The interaction between modulation-induced frequency chirp and chromatic dispersion (CD) is also among the main limiting factors in the implementation of DML-based IM/DD systems [2]. As throughput requirements increase beyond 100 Gbps per lane [2], these limitations become more challenging. Operating DMLs at symbol rates beyond 50 GBd, necessary for high data throughput, can introduce impairments such as nonlinear waveform distortion and more severe frequency chirping [14, 15, 16]. These impairments can reduce the effective signal-to-noise ratio (SNR) of the signal at the receiving end, limiting transmission distance and achievable data rates.
@@
Traditionally, digital signal processing (DSP) techniques for mitigating impairments in DML-based systems have focused on receiver-side equalization (EQ) or a combination of digital pre-distortion (DPD) at the transmitter and EQ at the receiver [17, 18, 2, 13, 19, 5]. These approaches optimize only one end of the system at a time and thus they may lead to suboptimal solutions compared to joint optimization techniques [20]. End-to-end (E2E) learning [21, 22, 23, 24] offers a promising approach to jointly optimizing the transmitter and receiver, potentially leading to optimal DSP configurations in terms of signal quality and power consumption. E2E learning in communication systems involves representing the transmitter and receiver as a single autoencoder (AE) artificial neural network (ANN) built around the transmission channel [25, 26]. This approach allows for the simultaneous optimization of the DSP at the transmitter (constellation shaping, pulse shaping, DPD) and receiver (EQ, symbol detection) while keeping the parameters of the channel model fixed.
@@
Applying E2E learning to DML-based systems is hindered by the physical model of the laser [27, 25, 28]. The DML large-signal dynamics are governed by nonlinear laser rate equations [29], which describe the interplay between carrier density and photon density in the cavity. These equations lack a closed-form large-signal solution, requiring the use of iterative numerical solvers to simulate the DML’s output. This simulation step is not directly differentiable, which prevents the backpropagation of gradients from the receiver to the transmitter and stalls the training process [30, 10]. Data-driven surrogate models based on machine learning offer a solution to the differentiability challenge posed by DMLs [24]. These models learn a differentiable representation of the DML dynamics based on input/output sequences. Data-driven models create a viable, differentiable pathway for backpropagation where analytical models fall short, unlocking the potential of end-to-end, gradient-based optimization for DML-based systems. Data-driven models can also be faster substitutes for computationally expensive ordinary differential equation (ODE) solvers [31] and derivative-free gradient approximators in parameter optimization [32]. Based on the gradients provided by the differentiable surrogate model, E2E learning allows the joint optimization of an arbitrarily large space of parameters that affect the performance of the system.
@@
In this paper, we demonstrate the surrogate-based E2E optimization of transmitter (TX) and receiver (RX) DSP and laser driving configurations over different lengths of standard single mode fiber (SSMF). The proposed data-driven surrogate model is trained to capture the chirp-CD interaction along the different fiber lengths, allowing the E2E optimization to mitigate its impact on performance. This is done by training a separate model for each transmission distance and baudrate. Our approach allows the E2E system to optimize the DSP (constellation shaping, pulse shaping, EQ and symbol detection) and laser-driving configurations (bias and modulation RF power) for every scenario. Thus, we extend our previous work in [33], showing the surrogate-based E2E optimization of TX and RX DSP and laser-driving configurations in a back-to-back (B2B) scenario. The proposed E2E system is benchmarked experimentally against two more traditional RX-side equalization based on root-raised cosine (RRC) pulse-shaped signals. The two RRC-shaped benchmark equalization strategies are based on (1) a linear (finite impulse response, FIR) and (2) a nonlinear (2nd order Volterra series, VNLE) equalizer to test the E2E system performance against both linear and nonlinear receiver-side equalization. We demonstrate, for the first time to our knowledge, the joint optimization of DSP and laser driving configurations in an experimental transmission setting. The E2E system achieves better error performance and eye opening than RX equalization across the studied fiber lengths. This improvement is realized while utilizing less bandwidth and employing lower modulation RF power and fewer filter taps than the RX-side equalization benchmarks.
@@ II DML characteristics
The characteristics of the laser play an important role in the error performance of DML-based IM/DD links, primarily due to the laser’s strong nonlinear dynamics, limited bandwidth, and significant frequency chirping. The utilized NEL NLK1551SSC DML is designed for 10 Gbps links for up to 50 km, featuring a typical nominal external efficiency of 0.15 W/A. The measured L-I curves of the utilized laser at three different controlled temperatures $T\in[22.5,25,27.5]$ °C are shown in Fig. 1a. The threshold current $I_{\mathrm{th}}$ of the laser is measured to be 7.5 mA. The three L-I curves show a similar pattern in their slope (differential external efficiency, $\eta_{d}$), where the value of $\eta_{d}$ is maximum around 52.5 mA (0.067 W/A) and decreases for both higher and lower $I_{\mathrm{bias}}$ values. The small-signal modulation responses measured for three different $I_{\mathrm{bias}}\in[50,70,90]$ mA and $T=25$ °C are plotted in Fig. 1b. The responses are measured by sweeping the frequency of an input sinusoidal electrical current and recording the corresponding magnitude of the modulated optical output power to characterize the bandwidth of the laser. The peak-to-peak current of the sinusoidal input signal is 8 mA. The measured 3-dB bandwidth for $I_{\mathrm{bias}}=[50,70,90]$ mA are $[14.2,17.1,18.7]$ GHz, respectively. This leads to the significant difference in the low frequency response of the 3 curves. Fig. 1 also hints at a trade-off between extinction ratio and modulation bandwidth (increasing with higher $I_{\mathrm{bias}}$ in Fig. 1b), with the $I_{\mathrm{bias}}\approx 50$ mA region offering a higher extinction ratio while delivering reduced bandwidth compared to higher bias currents.
@@ III Experimental modeling
Figure 2: Block diagram of the proposed experimental modeling approach. The digital domain is designed to build a data-driven model resembling the dynamics of the experimental setup (b). Acronyms and symbols:: 4PAM: 4-Pulse Amplitude Modulation; AWG: arbitrary waveform generator; $P_{\mathrm{RF}}$: modulation RF power; $G$: RF amplifier gain; $I_{\mathrm{bias}}$: bias current; DML: directly modulated laser; $L$: standard single-mode fiber length; PD: photodetector; DSO: digital storage oscilloscope; LSTM: Long-Short Term Memory; MSE: mean squared error.
Figure 4: Block diagram of the E2E scheme parameter optimization. The elements optimized by the scheme are highlighted in red. Acronyms and symbols:: 4PAM: 4-Pulse Amplitude Modulation; $P_{\mathrm{RF}}$: modulation RF power; $I_{\mathrm{bias}}$: bias current; LSTM: Long-Short Term Memory; MLSE: maximum likelihood sequence estimation.
To enable end-to-end optimization, we first train a differentiable data-driven model that captures the deterministic impairments in the link based on averaged experimental data. The averaging is performed to discard the stochastic components (noise) in the signal, simplifying the modeling task. The model imitates the static and dynamic response of the link from TX to RX, including laser dynamics, fiber-induced distortions, and photodetector nonlinearities. Given the memory effects and nonlinearity introduced by the DML, the structure of the data-driven model must include memory mechanisms and nonlinear activations that allow it to infer the response of the link accurately. Based on our previous work in DML modeling [31], we choose a long short term memory (LSTM) network structure for the task. LSTMs are recurrent neural networks (RNNs) aimed at the inference of sequential data with long temporal dependencies, making them ideal candidates to model the memory effects and nonlinearity associated with the response of DMLs. The core principle of LSTMs is the use of hidden states $\mathbf{h}_{t}$, that are updated for every time sample $t$ to retain information about the past inputs and outputs in order to provide temporal context. The balance between past and present information stored in $\mathbf{h}_{t}$ is determined through matrix multiplication and nonlinear activation, allowing LSTMs to operate as universal approximators [34]. Thus, the calculation of the output, or output gate $\mathbf{o}_{t}$ of an LSTM can be expressed as
@@
where $\mathbf{x}_{\mathrm{TX}}$ is the digital pulse-shaped signal to be transmitted, $N_{\mathrm{samps}}$ is the number of samples of $\mathbf{x}_{\mathrm{TX}}$. The model is trained on a dataset with varied waveforms, input powers ($P_{\mathrm{RF}}$), and bias currents ($I_{\mathrm{bias}}$) to ensure its robustness against fluctuations in these parameters. A 1024-sample pilot sequence is introduced before the 4PAM waveforms in order to allow simple symbol synchronization on the receiver side through autocorrelation. The experimental setup is organized into signal generation, transmission, and reception stages. The generation stage begins with a Keysight M8195A AWG (sampling rate: 65 GSa/s, bandwidth: 25 GHz, max $P_{\mathrm{RF}}$: 4 dBm). Its output is amplified by an SHF 806E RF amplifier ($G=13$ dB) and drives a NEL NLK1551SSC DML, which is temperature-stabilized at 25°C and biased using a DC power supply generating $I_{\mathrm{bias}}$. For transmission, the optical signal propagates through one of three different lengths ($L$) of SSMF: back-to-back, 1 km, or 2 km. At the receiver, a Coherent Corp. XPDV2020R photodetector (PD) with a nominal bandwidth of 50 GHz converts the signal. Finally, an Agilent DSOX93304Q oscilloscope (DSO sampling rate $R_{\mathrm{sa}}$: 80 GSa/s, bandwidth: 33 GHz) digitizes the received waveform experimentally. To match the transmitter hardware, the input sequences are linearly interpolated to 65 GSa/s before being sent to the AWG. However, the surrogate model is trained using linear interpolation to 80 GSa/s, matching the input and output sampling frequency to $R_{\mathrm{sa}}$. Once the waveforms have been propagated, the resulting distorted signal is captured 25 times and averaged in order to reduce stochastic effects, that cannot be captured by the LSTM model. The number of averages was chosen to balance sufficient noise suppression with a practical acquisition time. Once the full input/output dataset has been captured, averaged and synchronized, the data-driven LSTM model is trained using mean square error (MSE) between predicted and reference sequences as a loss function. The total size of the dataset is $10^{7}$ samples per baudrate and $L$ considered. The training is based on an Adam optimizer with learning rate $\mu_{\mathrm{model}}=1\times 10^{-4}$ and exponential decay rates $\beta_{1}=0.999,\beta_{2}=0.99$. The obtained surrogate model testing normalized mean square errors (NRMSEs) are listed in Table I. The surrogate model is able to deliver normalized mean errors around the 1% error mark throughout the $R_{s}$ and $L$ studied. However, the accuracy of the gradients provided by the surrogate model cannot be directly captured using MSE as a metric, and it must be assessed through the performance of the E2E optimization scheme.
@@ IV-A End-to-End scheme
where $\mathbf{n}\approx\mathbf{y}_{\mathrm{avg}}-\mathbf{y}_{\mathrm{raw}}$ is the estimated noise sequence, $\mathbf{y}_{\mathrm{avg}}$ is the output training sequence after averaging and $\mathbf{y}_{\mathrm{raw}}$ is one of the 25 sequences used during the averaging of $\mathbf{y}_{\mathrm{avg}}$. $P_{\mathbf{y}_{\mathrm{avg}}}$ and $P_{n}$ represent the average power of $\mathbf{y}_{\mathrm{avg}}$ and $\mathbf{n}$, respectively. Once $\mathrm{SNR}_{\mathrm{est}}$ is calculated, the variance of the AWGN noise is calculated as: $\sigma^{2}_{\mathrm{AWGN}}=P_{\mathbf{y}_{\mathrm{mod}}}/\mathrm{SNR}_{\mathrm{est}}$, where $\mathbf{y}_{\mathrm{mod}}$ is the output sequence of the surrogate model. The noisy sequences predicted by the surrogate model are then processed by the RX DSP. The received 80 GSa/s sequences are linearly downsampled to 2 SpS before equalization. The receiver-side equalization is based on a trainable 15 $T/2$-spaced-tap FFE. The equalized signal is then downsampled to 2 SpS and detected using 1-tap maximum likelihood sequence estimation (MLSE) based on a trainable feedforward neural network (FFNN). The MLSE FFNN is based on a $1\times 32$ ReLU-activated feedforward input layer, followed by a hidden $32\times 32$ feedforward layer with identical activation. The output $32\times 4$ feedforward layer is activated by a softmax function. The MLSE FFNN serves as an approximation to the non-continuous, non differentiable thresholding function, delivering symbol probabilities as an output (soft decision). The link optimization uses categorical cross-entropy $H$ as loss function:
@@
Once the optimization phase is complete, the E2E scheme is validated on the experimental setup to verify its performance. This is done by applying the DSP, $P_{\mathrm{RF}}$ and $I_{\mathrm{bias}}$ configurations obtained in the offline optimization to the experimental setup. Given that the optimized $x_{\mathrm{TX}}$} represents voltage levels, the only change applied to the TX DSP on the experimental setup is the downsampling of $x_{\mathrm{TX}}$} from 80 to 65 GSa/s, in order to match the sampling frequency of the AWG. The metric utilized to measure the performance of the E2E scheme is SER. The SER calculation on the E2E scheme is performed based on $a_{k}$ and $r_{k}$, following:

### v2

@@ I Introduction
The explosive growth of internet services and data traffic, driven by the surge in cloud computing and artificial intelligence (AI) applications, is placing considerable pressure on the global communications infrastructure [10273773]. Specifically, short-reach intra-data center interconnects must continue to increase data throughput while maintaining low costs. Intensity modulation and direct detection (IM/DD) is the dominating technology for such interconnects [10238466, app13148125, 10526441]. Directly modulated lasers (DMLs) are an attractive technology for short-reach IM/DD optical links, offering low energy consumption, compact size, and affordability [10528667, 10438718, 10080935, 10323176, 10041217, 10261204]. This makes them well-suited for communication systems demanding high data rates at minimal cost. DMLs are simpler and potentially more efficient than alternative transmitter technologies [10041217]. However, the limited modulation bandwidth of DMLs can restrict their achievable symbol rates [Sudo:21, 9874980, 8570831]. The interaction between modulation-induced frequency chirp and chromatic dispersion (CD) is also among the main limiting factors in the implementation of DML-based IM/DD systems [10238466]. As throughput requirements increase beyond 100 Gbps per lane [10238466], these limitations become more challenging. Operating DMLs at symbol rates beyond 50 GBd, necessary for high data throughput, can introduce impairments such as nonlinear waveform distortion and more severe frequency chirping [575810, 10925524, Zhang:25]. These impairments can reduce the effective signal-to-noise ratio (SNR) of the signal at the receiving end, limiting transmission distance and achievable data rates.
@@
Traditionally, digital signal processing (DSP) techniques for mitigating impairments in DML-based systems have focused on receiver-side equalization (EQ) or a combination of digital pre-distortion (DPD) at the transmitter and EQ at the receiver [Kottke:17, Xu:20, 10238466, 8570831, 8259239, 10528667]. These approaches optimize only one end of the system at a time and thus they may lead to suboptimal solutions compared to joint optimization techniques [10839011]. End-to-end (E2E) learning [8054694, Karanov:20, 10824218, Hernandez:24] offers a promising approach to jointly optimizing the transmitter and receiver, potentially leading to optimal DSP configurations in terms of signal quality and power consumption. E2E learning in communication systems involves representing the transmitter and receiver as a single autoencoder (AE) artificial neural network (ANN) built around the transmission channel [10093964, 10124361]. This approach allows for the simultaneous optimization of the DSP at the transmitter (constellation shaping, pulse shaping, DPD) and receiver (EQ, symbol detection) while keeping the parameters of the channel model fixed.
@@
Applying E2E learning to DML-based systems is hindered by the physical model of the laser [Agrell_2024, 10093964, 10530894]. The DML large-signal dynamics are governed by nonlinear laser rate equations [7972948], which describe the interplay between carrier density and photon density in the cavity. These equations lack a closed-form large-signal solution, requiring the use of iterative numerical solvers to simulate the DML’s output. This simulation step is not directly differentiable, which prevents the backpropagation of gradients from the receiver to the transmitter and stalls the training process [10059125, 10261204]. Data-driven surrogate models based on machine learning offer a solution to the differentiability challenge posed by DMLs [Hernandez:24]. These models learn a differentiable representation of the DML dynamics based on input/output sequences. Data-driven models create a viable, differentiable pathway for backpropagation where analytical models fall short, unlocking the potential of end-to-end, gradient-based optimization for DML-based systems. Data-driven models can also be faster substitutes for computationally expensive ordinary differential equation (ODE) solvers [10382548] and derivative-free gradient approximators in parameter optimization [9513589]. Based on the gradients provided by the differentiable surrogate model, E2E learning allows the joint optimization of an arbitrarily large space of parameters that affect the performance of the system.
@@
In this paper, we demonstrate the surrogate-based E2E optimization of transmitter (TX) and receiver (RX) DSP and laser driving configurations over different lengths of standard single mode fiber (SSMF). The proposed data-driven surrogate model is trained to capture the chirp-CD interaction along the different fiber lengths, allowing the E2E optimization to mitigate its impact on performance. This is done by training a separate model for each transmission distance and baudrate. Our approach allows the E2E system to optimize the DSP (constellation shaping, pulse shaping, EQ and symbol detection) and laser-driving configurations (bias and modulation RF power) for every scenario. Thus, we extend our previous work in [Hernandez:25] in three different ways. Firstly, while [Hernandez:25] only addressed a back-to-back system, here we also study the performance of the system under transmission through 1 and 2 km of standard single-mode fiber (SSMF). This additional analysis assesses the system’s capability to mitigate undesirable effects from the interaction between chromatic dispersion and frequency chirp. Second, we bypass the digital bandwidth limitation present in our previous work, simplifying the structure of the surrogate model. Finally, we analyze and interpret the TX and RX filters obtained through numerical optimization. The analysis clarifies the trade-offs taken by the E2E scheme and explains the results achieved.Although the achieved throughput does not reach that of state-of-the-art DMLs due to device limitations, the presented results aim to validate the proposed methods in terms of the potential performance improvements. The proposed E2E system is benchmarked experimentally against two more traditional RX-side equalization based on root-raised cosine (RRC) pulse-shaped signals. The two RRC-shaped benchmark equalization strategies are based on (1) a linear (finite impulse response, FIR) and (2) a nonlinear (2nd order Volterra series, VNLE) equalizer to test the E2E system performance against both linear and nonlinear receiver-side equalization. We demonstrate, for the first time to our knowledge, the joint optimization of DSP and laser driving configurations in an experimental transmission setting. The E2E system achieves better error performance and eye opening than RX equalization across the studied fiber lengths. This improvement is realized while utilizing less bandwidth and employing lower modulation RF power and fewer filter taps than the RX-side equalization benchmarks.
@@ II DML characteristics
The characteristics of the laser play an important role in the error performance of DML-based IM/DD links, primarily due to the laser’s strong nonlinear dynamics, limited bandwidth, and significant frequency chirping. The utilized NEL NLK1551SSC DML is designed for 10 Gbps links for up to 50 km, featuring a typical nominal external efficiency of 0.15 W/A. The bias current, or direct current component of the signal to the laser $I_{\mathrm{bias}}$ is limited to a maximum of 100 mA. The measured L-I curves of the utilized laser at three different controlled temperatures $T\in[22.5,25,27.5]$ °C are shown in Fig. 1a. The threshold current $I_{\mathrm{th}}$ of the laser is measured to be 7.5 mA. The three L-I curves show a similar pattern in their slope (differential external efficiency, $\eta_{d}$), where the value of $\eta_{d}$ is maximum around 52.5 mA (0.067 W/A) and decreases for both higher and lower $I_{\mathrm{bias}}$ values. The small-signal modulation responses measured for three different $I_{\mathrm{bias}}\in[50,70,90]$ mA and $T=25$ °C are plotted in Fig. 1b. The responses are measured by sweeping the frequency of an input sinusoidal electrical current and recording the corresponding magnitude of the modulated optical output power to characterize the bandwidth of the laser. The peak-to-peak current of the sinusoidal input signal is 8 mA. The measured 3-dB bandwidth for $I_{\mathrm{bias}}=[50,70,90]$ mA are $[14.2,17.1,18.7]$ GHz, respectively.
This leads to the significant difference in the low frequency response of the 3 curves. Fig. 1 also hints at a trade-off between extinction ratio and modulation bandwidth (increasing with higher $I_{\mathrm{bias}}$ in Fig. 1b), with the $I_{\mathrm{bias}}\approx 50$ mA region offering a higher extinction ratio while delivering reduced bandwidth compared to higher bias currents.
@@ III Experimental modeling
Figure 3: Block diagram of the proposed experimental modeling approach. The digital domain (a) is designed to build a data-driven model resembling the dynamics of the experimental setup (b). Acronyms and symbols: 4PAM: 4-Pulse Amplitude Modulation; AWG: arbitrary waveform generator; $\mathbf{P_{\mathrm{RF}}}$: modulation RF power; $\mathbf{G}$: RF amplifier gain; $\mathbf{I_{\mathrm{bias}}}$: bias current; DML: directly modulated laser; $\mathbf{L}$: standard single-mode fiber length; PD: photodetector; DSO: digital storage oscilloscope; LSTM: Long-Short Term Memory; MSE: mean squared error.
To enable end-to-end optimization, we first train a differentiable data-driven model that captures the deterministic impairments in the link based on averaged experimental data. The model is intended as part of the offline calibration of the transmission system, and it does not need to run in real time. Therefore, averaging is performed to discard the stochastic components (noise) in the signal, simplifying the modeling task. The model imitates the static and dynamic response of the link from TX to RX, including laser dynamics, fiber-induced distortions, and photodetector nonlinearities. Given the memory effects and nonlinearity introduced by the DML, the structure of the data-driven model must include memory mechanisms and nonlinear activations that allow it to infer the response of the link accurately. Based on our previous work in DML modeling [10382548], we choose a long short term memory (LSTM) network structure for the task. LSTMs are recurrent neural networks (RNNs) aimed at the inference of sequential data with long temporal dependencies, making them ideal candidates to model the memory effects and nonlinearity associated with the response of DMLs. The core principle of LSTMs is the use of hidden states $\mathbf{h}_{t}$, that are updated for every time sample $t$ to retain information about the past inputs and outputs in order to provide temporal context. The balance between past and present information stored in $\mathbf{h}_{t}$ is determined through matrix multiplication and nonlinear activation, allowing LSTMs to operate as universal approximators [SIEGELMANN1995132]. Thus, the calculation of the output, or output gate $\mathbf{o}_{t}$ of an LSTM can be expressed as
@@
where $\mathbf{x}_{\mathrm{TX}}$ is the digital pulse-shaped signal to be transmitted, $N_{\mathrm{samps}}$ is the number of samples of $\mathbf{x}_{\mathrm{TX}}$. The model is trained on a dataset with varied waveforms, input powers ($P_{\mathrm{RF}}$), and bias currents ($I_{\mathrm{bias}}$) to ensure its robustness against fluctuations in these parameters. A 1024-sample pilot sequence is introduced before the 4PAM waveforms in order to allow simple symbol synchronization on the receiver side through autocorrelation. The experimental setup is organized into signal generation, transmission, and reception stages. The generation stage begins with a Keysight M8195A AWG (sampling rate: 65 GSa/s, bandwidth: 25 GHz, max $P_{\mathrm{RF}}$: 4 dBm). Its output is amplified by an SHF 806E RF amplifier ($G=13$ dB, bandwidth: 38 GHz) and drives a NEL NLK1551SSC DML, which is temperature-stabilized at 25°C and biased using a DC power supply generating $I_{\mathrm{bias}}$. For transmission, the optical signal propagates through one of three different lengths ($L$) of SSMF: back-to-back, 1 km, or 2 km. At the receiver, a Coherent Corp. XPDV2020R photodetector (PD) with a nominal bandwidth of 50 GHz and a nominal responsivity of 0.65 A/W converts the signal. Finally, an Agilent DSOX93304Q oscilloscope (DSO sampling rate $R_{\mathrm{sa}}$: 80 GSa/s, bandwidth: 33 GHz) digitizes the received waveform experimentally. To match the transmitter hardware, the input sequences are linearly interpolated to 65 GSa/s before being sent to the AWG. However, the surrogate model is trained using linear interpolation to 80 GSa/s, matching the input and output sampling frequency to $R_{\mathrm{sa}}$.
Figure 4: Block diagram of the E2E scheme parameter optimization. The elements optimized by the scheme are highlighted in red. Acronyms and symbols: 4PAM: 4-Pulse Amplitude Modulation; $\mathbf{P_{\mathrm{RF}}}$: modulation RF power; $\mathbf{I_{\mathrm{bias}}}$: bias current; LSTM: Long-Short Term Memory; MLSE: maximum likelihood sequence estimation.
@@
Once the waveforms have been propagated, the resulting distorted signal is captured 25 times and averaged in order to reduce stochastic effects, that cannot be captured by the LSTM model. The number of averages was chosen to balance sufficient noise suppression with a practical acquisition time. Once the full input/output dataset has been captured, averaged and synchronized, the data-driven LSTM model is trained using mean square error (MSE) between predicted and reference sequences as a loss function. The total size of the dataset is $10^{7}$ samples per baudrate and $L$ considered. Although training one model per baudrate and $L$ may increase operational complexity, it yields better modeling accuracy for each studied scenario. The training is based on an Adam optimizer with learning rate $\mu_{\mathrm{model}}=1\times 10^{-4}$ and exponential decay rates $\beta_{1}=0.999,\beta_{2}=0.99$. The obtained surrogate model testing normalized mean square errors (NRMSEs) are listed in Table I. The surrogate model is able to deliver normalized mean errors around the 1% error mark throughout the $R_{s}$ and $L$ studied. However, the accuracy of the gradients provided by the surrogate model cannot be directly captured using MSE as a metric, and it must be assessed through the performance of the E2E optimization scheme.
@@ IV-A End-to-End scheme
where $\mathbf{n}\approx\mathbf{y}_{\mathrm{avg}}-\mathbf{y}_{\mathrm{raw}}$ is the estimated noise sequence, $\mathbf{y}_{\mathrm{avg}}$ is the output training sequence after averaging and $\mathbf{y}_{\mathrm{raw}}$ is one of the 25 sequences used during the averaging of $\mathbf{y}_{\mathrm{avg}}$. $P_{\mathbf{y}_{\mathrm{avg}}}$ and $P_{n}$ represent the average power of $\mathbf{y}_{\mathrm{avg}}$ and $\mathbf{n}$, respectively. Once $\mathrm{SNR}_{\mathrm{est}}$ is calculated, the variance of the AWGN noise is calculated as: $\sigma^{2}_{\mathrm{AWGN}}=P_{\mathbf{y}_{\mathrm{mod}}}/\mathrm{SNR}_{\mathrm{est}}$, where $\mathbf{y}_{\mathrm{mod}}$ is the output sequence of the surrogate model. The noisy sequences predicted by the surrogate model are then processed by the RX DSP. The received 80 GSa/s sequences are linearly downsampled to 2 SpS before equalization. The receiver-side equalization is based on a trainable 15 $T/2$-spaced-tap FFE. The equalized signal is then downsampled to 2 SpS and detected using 1-tap maximum likelihood sequence estimation (MLSE) based on the trainable feedforward neural network (FFNN) shown in Fig. 5. The MLSE FFNN is based on a $1\times 32$ ReLU-activated feedforward input layer, followed by a hidden $32\times 32$ feedforward layer with identical activation. The output $32\times 4$ feedforward layer is activated by a softmax function. The MLSE FFNN serves as an approximation to the non-continuous, non differentiable thresholding function, delivering symbol probabilities as an output (soft decision). The link optimization uses categorical cross-entropy $H$ as loss function:
@@
Once the optimization phase is complete, the E2E scheme is validated on the experimental setup to verify its performance. This is done by applying the DSP, $P_{\mathrm{RF}}$ and $I_{\mathrm{bias}}$ configurations obtained in the offline optimization to the experimental setup. Given that the optimized $x_{\mathrm{TX}}$} represents voltage levels, the only change applied to the TX DSP on the experimental setup is the downsampling of $x_{\mathrm{TX}}$} from 80 to 65 GSa/s, in order to match the sampling frequency of the AWG. The metric utilized to measure the performance of the E2E scheme is SER. This is due to the per-symbol nature of the proposed E2E optimization scheme, which allows evaluating signal quality regardless of the bit mapping employed. The SER calculation on the E2E scheme is performed based on $a_{k}$ and $r_{k}$, following: